A man buys milk at a certain rate per liter and after adding water to it sells the mixture at the same rate. Find in what ratio he much mix water to milk so as to gain 20% on his outlay?

Answer by ptaylor(2198)

Let x=rate per liter that he pays for milk and also the rate that he sells the milk mixture
Let y=amount of liters of milk initially purchased
And let w=amount of water that needs to be added
Cost of milk =xy
Value of the milk mixture=x(y+w)
A 20% gain in outlay can be written as follows: (% gain =(new amt-old amt)/old amt)
(x(y+w)-xy)/xy=0.2
Notice that the cost is not relevant
((y+w)-y)/y=0.2 or
w/y=0.2=1/5
This tells us that for every six liters of mixture, 5 liters is milk and 1 liter is water
Hope this helps---ptaylor